1.5gm of a monobasic acid when disslved in 150g of water lowers Frezing pt by 0.165 degrees celsius.


0.5g of same acid when titrated,after dissolution in water, requires 37.5ml of N/10 alkali.Calculate degree of dissociation of acid(If Kf of water is 1.86)

Grade:12

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Freezing point depression can be calculated using the Clausius-Clapeyron equation and Raoult's law. In a dilute ideal solution the freezing point is:

Freezing Pointtotal = Freezing Pointsolvent - ΔTf

where ΔTf = molality * Kf * i

Kf = cryoscopic constant (1.86°C kg/mol for the freezing point of water)

i = Van't Hoff factor

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sourabh Dhavala
10 years ago

The answer is 18.34%. I''m not getting that answer. 

tanay
5 years ago
∆Tb=(W2 *1000)/(Mw2  × W1)
0.168=(1.86 X 1000)/(M X 100)
M = 11 0.71
 
N1 xV1   = Nx V
15.1/10 =(0.2×1000)/M.M   (equivalent)
= 132.45
i= 132.45× 110.71 = 1+ a            HX= H+ X-
alpha= 19.6